Hi,
I been trying to solve this problem but I can't find the right formula to do it. I tried A-1/2b*h but I don't get the answer right. Would you help me please?
Thank you
the face of the triangular concrete panel has an area of 45 square meters and its base is 3 meters longer than twice its height. find the length of its base

Answer by TimothyLamb(4379)

a = (1/2)bh
---
b = 2h + 3
---
45 = (1/2)bh
45 = (1/2)(2h + 3)h
45 = (1/2)(2hh + 3h)
45 = hh + (3/2)h
a(h) = hh + (3/2)h - 45
---
the above quadratic equation is in standard form, with a=1, b=1.5, and c=-45
---
to solve the quadratic equation, plug this:
1 1.5 -45
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the roots of the quadratic are:
6
-7.5
---
the negative root doesn't make sense as length, so use the positive root:
---
Answer:
h = 6 m
b = 15 m
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php